OpenStudy (alizandraw):

According to the equation 2Na + 2H2O mc012-1.jpg 2NaOH+H2, what mass of Na is required to yield 22.4 L of H2 at STP? (The atomic mass of Na is 22.99 u.) 1.00 g 2.00 g 23.0 g 46.0 g

OpenStudy (abmon98):

2Na+2H2O-->2NaOH+H2 at STP The standard temperature is 273 K (0° Celsius) and the standard pressure is 1 atm pressure. At STP, one mole of gas occupies 22.4 L of volume (molar volume). so number of moles of H2 is 1 look at the mole ratio of sodium to hydrogen its 2:1

OpenStudy (abmon98):

2:1 x:1 x=1*2/1=2 moles of sodium Number of moles=Mass(g)/Molar Mass(g/mol)